Grade Upto college levelGeneral Physics

A cell of emf 1.5V is connected with an ammeter of resistance 0.05ohm.If the current in the circuit is 2A the internal resistance of the cell is...???

Srijan
12 Years agoGrade Upto college level

1 Answer

Bhausaheb
7 Years ago
Formula for internal resistance of cell is
r = (E - V) /I
Or
r = (E - IR) /I
Given quantities in problem
E = 1.5 V
R = 0.05 ohm
and I = 2 amp
r = (1.5 - 2*0.05)/2
r = 0.7 ohm